Raymond Mining Corporation has 8.5 million shares of common stock outstanding and 135,000

semi-annual bonds outstanding with face value $1000 and coupon rate of 7.5%. The common

stock currently sells for $34 per share and has a beta of 1.25, and the bonds have 15 years to

maturity and sell for 114 percent of the face value ($114 per $100 of face value). The market risk

premium is 7.5 percent, T-bills are yielding 4 percent (thus risk free rate is 4%), and the Raymond

Mining tax rate is 35%. Raymond Mining is considering an investment which will cost $2,590,000.

The investment produces no cash flows for the first year. In the second year, the cash inflow is

$580,000. This inflow will increase to $1,500,000 and then $2,000,000 for the following years before

ceasing permanently. What is the project’s NPV? What is the Internal Rate of Return? What is the

Discounted Payback Period? What is the Profitability Index?
Answer #1

Market value of equity, E = P x N = 34 x 8.5 million = 289,000,000

Market value of Debt, D = P x N = 114% x 1,000 x 135,000 = $  153,900,000

Ke = Rf + Beta x RMp = 4% + 1.25 x 7.5% = 13.375%

Kd = 2 x RATE (Nper, PMT, PV, FV) = 2 x RATE (2 x 15, 7.5% / 2 x 1000, -114% x 1000, 1000) = 6.065%

Proportion of debt in capital structure = Wd = D / (D + E) = 153,900,000 / (153,900,000 + 289,000,000) = 34.75%

Proportion of equity, We = 1 - Wd = 1 - 34.75% = 65.25%

Tax rate, T = 35%

Hence, discount rate = WACC = r = Wd x Kd x (1 - T) + We x Ke = 10.097%

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $. Adjacent cells in blue contain the formula in excel I have used to get the final output.